# SNIC: Synthesized Noisy Images using Calibration

### v1

@@
SNIC: Synthesized Noisy Images using Calibration
Nik Bhatt
Columbia University
nb3097@columbia.edu
1 Abstract
Advanced denoising algorithms require large, high-quality datasets. Physics-based, statistical noise models can create such datasets by realistically simulating noise in digital images. However, there is little information on the correct way to calibrate and tune these heteroscedastic models, and a lack of published datasets using them. In this paper, we explore the process of building high-quality heteroscedastic noise models. Our methods produce realistic synthesized noisy images, in both RAW and TIFF formats. Our synthesized noisy images achieve comparable LPIPS results to real noisy images, and greatly outperform those created with manufacturer-provided DNG noise models both in LPIPS and when tested with a state-of-the-art (SOTA) denoising model. Using our approach, we created the Synthesized Noisy Images using Calibration dataset (SNIC) containing over 6000 noisy images, comprising 30 scenes from four sensors, including two smartphone sensors, a point-and-shoot, and a DSLR. SNIC is the first synthesized noisy image dataset provided in both RAW and TIFF format.
@@ 2 Introduction
All digital images contain noise as a by-product of the capture process. Noise arises from the quantum effects of photon arrival, and from various electronic components in the sensor and camera circuitry, including sensor readout, amplifiers, and analog-to-digital conversion. Images captured in low-light environments, and especially with the smaller sensors found on smartphones, typically include considerable amounts of noise, negatively affecting the appearance and aesthetics of the images. Thus, noise reduction, or denoising, is an important and difficult task that has been the focus of active research for years. Denoising is an ill-posed inverse problem with no one correct solution, but rather an infinite set of possible solutions. Denoising is best performed on RAW images (rather than camera-generated JPEGs) because RAW images are scene-referred, undergo less processing by the image signal processor (ISP), contain linear data, and do not include gamma correction.
Currently most research in denoising focuses on machine learning models. Thus, researchers developing new denoising algorithms rely on large image datasets to train and test their models. There are three main approaches to acquiring an adequate dataset.
@@
  • Second, one can use an existing real noisy dataset that consists of low ISO and high ISO images that have been extremely carefully aligned. Making a real noise dataset is exceptionally difficult, so there are only a few available. Most were created in 2017 or 2018, and thus use older cameras such as the iPhone 7, released over nine years ago. Real noise datasets are also often limited to specially lit indoor scenes.
  • Third, one can add noise to one’s own images using a heteroscedastic Poisson-Gaussian noise model, which is superior to a homoscedastic one like AWGN. A heteroscedastic model can be easily retrieved from DNG images, though we found those models are often inaccurate and incomplete. We have chosen a variant of the third approach: create a custom physics-based heteroscedastic noise model by calibrating cameras and then tuning the model with validation images. This requires more effort than simply using DNG noise profiles, but our models provide perfectly aligned, high-quality noisy RAW images. Further, this approach offers full control over the choice of cameras and scenes used.
The choice and quality of the dataset directly affects the performance of a denoising algorithm on real images. For physically modeled datasets, the key is minimizing the discrepancy between the synthetic noise and actual noise [1]. In this paper, we focus on physics-based noise modeling, using careful calibration to determine photon shot noise, combined with complete capture of signal-independent noise sources via dark frames. We confirm the inaccuracy of manufacturer DNG noise profiles. We demonstrate how to create and tune a heteroscedastic, physically-based noise model to produce accurate synthesized noise, and use our models to create a large synthesized noisy image dataset, with both RAW and TIFF images. Our noisy images produce excellent results both in LPIPS measurements and in denoising tests.
@@
  • Heteroscedastic noise models that achieve a 54% to 64% improvement in PSNR dB vs. standard DNG noise models.
@@ 3 Related Work
In this section, we will discuss the principal ways in which noise models and datasets are created to train machine learning denoising models. While learning-based noise generation methods exist [2], this paper focuses on physics-based statistical approaches for dataset creation, and we do not discuss denoising models in depth, of which there are many.
@@ 3.1 Additive White Gaussian Noise
Additive White Gaussian Noise (AWGN) is a simple noise model. AWGN adds noise to all pixels, regardless of intensity, using a simple Gaussian distribution with zero mean and a chosen variance [2]:
@@ 3.2 Poisson-Gaussian Noise (P-G)
Poisson-Gaussian (P-G) is a physics-based noise model which improves upon AWGN by including both a signal-dependent component (shot noise), and a signal-independent component (read noise). Shot noise occurs because photon arrival is a quantum random process. For a pixel with mean signal $\mu$, the actual observed photon count includes random photon shot noise $N_{p}$. The total photon count follows a Poisson distribution [1]:
@@ 3.2.1 DNG Noise Profiles
Adobe, Inc. is the creator of the Digital Negative Specification (DNG) which is an open format for encoding digital images [3]. DNG files can store either RAW or non-RAW images. The DNG NoiseProfile tag (51041) uses the Poisson-Gaussian formulation. There are some weaknesses in the NoiseProfile specification. First, the specification simply contains one or more pairs of numbers (the scaling and bias parameters from Poisson-Gaussian). It does not include the ISO for which the noise applies, and there is no way to provide the parameters for multiple ISO values. Second, while some DNG NoiseProfiles include the parameters for each of the Bayer pattern channels (R, G1, G2, and B), it is very common to find NoiseProfiles that contain only three pairs of numbers (R, G, B), or simply a single pair of numbers which are then applied to all four channels equally (iPhone DNGs contain only this single pair). Since the ISO is not specified in the NoiseProfile tag, one must examine DNG files at different ISOs to determine the actual DNG model.
Finally, beyond the lack of per-channel information, DNG noise profiles provided by camera manufacturers tend to be rather inaccurate, compared to calibrated profiles [1]. We have found that profiles generated by Adobe for proprietary RAW files (such as the Sony ARW format) can be more accurate, but still retain the other issues described above.
@@ 3.2.2 Calibration and Dark Current
Modeling noise accurately using a Poisson-Gaussian formulation requires careful calibration (described in Methods). In 1994, [4] discusses calibrating a CCD using a gray card, as well as extracting Fixed Pattern Noise by shifting both the light source and gray card. In our calibrations, we confirmed the finding by [1] that fixed pattern noise is not a relevant component in modern CMOS sensors. Thus, our method for calibration deviates somewhat from [4] and includes a tuning portion using real noisy validation images at a range of ISO levels.
The signal-independent part of noise appears in all images from a given sensor, even completely dark ones (e.g., with a lens cap on). Any electronics component can produce noise [1]; as a result, it is extremely difficult, if not impossible, to fully characterize and model this noise physically. From [1] is the complete formulation for a digital image D:
@@
where $K_{a}$ and $K_{d}$ are the analog and digital gains, $N_{1}$ includes noise affected by analog gain, $N_{2}$ includes thermal and column fixed pattern noise that takes place after the analog gain step, and $N_{q}$ is quantization noise when converting analog values to digital ones.
There are a few techniques for capturing signal-independent noise for the purpose of noise synthesis. The simplest (and most common) is to compute it as part of the linear regression required for P-G noise modeling. The second is to capture dark frames (frames with no light hitting the sensor, such as with a lens cap, and/or in a completely dark room). This second approach, described by [1] as dark current, is used in lieu of the Poisson-Gaussian bias term. It more accurately captures sensor independent noise, as dark images include all noise sources by definition. However, as different sensors (and camera electronics) produce different amounts of noise (and at different stages), we have found that the value of capturing dark current varies substantially from sensor to sensor. Specifically, professional-quality cameras can be very accurately modeled using a standard P-G linear regression. Further, some image signal processors (ISPs) in cameras, such as the iPhone 11 Pro, will process RAW images to suppress or alter noise based on the contents of the scene.
@@ 3.3 Denoising Datasets
There are a number of denoising datasets, falling into two categories: ones that use AWGN to synthesize noise, and ones that use real noisy images. Due to the quality issues with AWGN, and the lack of published P-G datasets, we only list real noise datasets. Table 1 lists some popular real noisy image datasets [5], [6], [7], [8], [9], [10], [11].
@@
Table 1: Real Noise Datasets. Note: Most RAW datasets listed are from 2017-2018, highlighting the need for more updated datasets.
Real noise datasets group low-ISO clean images with noisy images captured at various ISO values [9]. Perfect alignment is challenging; even sub-millimeter camera movement can cause misalignment. Further, strict lighting requirements limit most datasets to indoor scenes. For example, SIDD contains 30,000 images but only 10 distinct scenes [9], while DND provides just 50 image pairs [8]. Most real noise datasets were created in 2017-2018, predating recent sensor advances. Perfectly aligned real noise is a high bar, but when achieved, it provides an unimpeachable ground truth for training. Some datasets, including [10] and [11], only include JPEG or PNG images, so they are not usable for RAW-based denoising.
Aside from AWGN, the alternative to real noise is physically modeled, Poisson-Gaussian synthesis. DNG RAW images include P-G noise profiles provided by the manufacturer or by Adobe as part of its camera calibration process. Because the quality of P-G models varies greatly, we created our own high-quality noise models. While somewhat laborious to create, our models have advantages over real noise. Our noisy images are perfectly aligned by definition, both indoor and outdoor scenes can be captured with a variety of lighting, and there is no limit to the number of images that can be produced. Further, by using dark current frames for some sensors, complex noise sources can be modeled.
@@ 4.1 RAW Images and Noise
Most cameras (including the ones we used) use a CMOS chip with a color filter array placed over the sensor sites. The color filter array usually follows the Bayer pattern (also called a mosaic), which arranges red, green, and blue filters in a prescribed order. There are twice as many green pixels as either red or blue pixels, due to the human eye’s increased sensitivity to green. There are two principal ways to process images for noise modeling and synthesis. One is to demosaic the Bayer data into RGB data and measure that data (or simply use TIFF, PNG, or JPEG data). The second is to operate on the Bayer data directly. Our approach works with the Bayer data directly because it provides higher quality. Bayer data is linear, rather than gamma-corrected, and we can ignore white balance and the specifics of the demosaic algorithm or the tone curve applied (of which there are many). Further, the RAW-based noise model can produce different values for each of the two green channels which is not possible once they have been processed into RGB.
@@ 4.2 DNG Noise Profile Analysis
The Poisson-Gaussian formulation enables straightforward synthesis of noisy images. However, the quality of this synthesis is entirely dependent on the accuracy and completeness of the formulation. One can either build a noise model manually, or rely on published noise profiles. Because building profiles manually can be labor-intensive and error-prone, most researchers rely on profiles stored as DNG NoiseProfile tags. For proprietary RAW files, the simplest approach is to convert RAW files into DNGs and then read the NoiseProfile built by Adobe’s DNG Converter software. This noise profile is computed by Adobe as part of its process for supporting proprietary RAW files. As mentioned previously, the DNG NoiseProfile specification does not include ISO data, so to build an accurate P-G model, one must provide the DNG converter a series of RAW files, shot with the desired range of ISO values. We found that for two Sony cameras (RX100 IV and A7R III), the Adobe profiles are very similar to the ones we built through our calibration and include noise parameters for red, green, and blue channels (though not for both green channels).
For cameras (including smartphone cameras) that produce DNG files natively, the noise profiles are provided by the manufacturer, not Adobe. We found that the DNG noise profile for the iPhone 11 Pro is both incomplete and inaccurate, especially at higher ISO values, which confirms the findings of [1]. First, the DNG NoiseProfile for the iPhone 11 Pro does not include scaling and offset values for each channel, providing instead only a single pair of values. Second, as the ISO rises, the scaling factor should rise, since noise is related to analog gain (ISO). However, for the iPhone 11 Pro, the values level off. This is discussed further in the next section.
@@
Figure 3: DNG NoiseProfile scaling term (a) for iPhone 11 Pro and iPhone 15 Pro across ISO values. Both models show non-monotonic behavior at high ISOs.
@@ 4.3 Non-linearity in iPhone Noise Models
We observed a distinct difference between the noise models for the Sony cameras and the iPhone main and telephoto cameras. While the Sonys had a fairly consistent linear relationship (in log-log space) with ISO, the models for the iPhone sensors sharply level off, which is physically impossible. The Sony noise models were created by Adobe, while the iPhone captures in DNG format already, so the models are written into the original files by Apple. This leveling-off would point to one of two explanations: either the calibration was performed incorrectly, or the images used to compute the noise parameters do not show the expected noise. Because we replicated this same leveling-off in our own calibration process for the iPhone cameras, it pointed to the latter explanation. Visually inspecting RAW DNG images confirmed our suspicions: at the same ISO and exposure time, despite being Bayer pattern RAW images, our flat-field iPhone images do not have the same noise characteristics as validation RAW images containing texture and complexity.
@@
Validation Image
Figure 4: Crops of calibration and validation iPhone 11 Pro images shot at the same ISO and exposure times.
Thus, to produce iPhone images with accurately synthesized noise, we found our calibrated model must be further tuned using carefully captured validation images at a range of ISO values.
We validated our models using 6 scenes captured across the four cameras (3 scenes each captured with the iPhone and Sony RX100 IV, and 3 additional scenes captured with the Sony A7R III). Each validation scene was captured at a low ISO (clean reference) and at 6-7 noisy ISO levels in powers of two, with two noisy captures per ISO level for baseline noise analysis. This yielded 156 full-resolution RAW validation images spanning most or all of the true ISO range of each camera. All validation images were captured using the same tripod and tethering setup as calibration to minimize motion. Unlike the flat-field calibration images, the validation images contained diverse textures, colors, and content, with a mix of indoor and outdoor scenes.

### v2

@@
Abstract
Advanced denoising algorithms require large, high-quality datasets. Physically-based statistical noise models can create such datasets by realistically simulating noise in digital images. However, there is little information on the correct way to calibrate and tune these heteroscedastic models, and a lack of published datasets using them. In this paper, we explore the process of building high-quality heteroscedastic noise models. Our methods produce realistic synthesized noisy images in both RAW and TIFF formats. Our synthesized noisy images achieve comparable LPIPS results to real noisy images; when tested with a state-of-the-art denoising model, our images reduce the PSNR gap versus real noise by 54–64% compared to those synthesized using manufacturer-provided DNG noise models. Using our approach, we created the Synthesized Noisy Images using Calibration dataset (SNIC) containing over 6000 noisy images, comprising 30 scenes from four sensors, including two smartphone sensors, a point-and-shoot, and a DSLR. SNIC is the first synthesized noisy image dataset provided in both RAW and TIFF format.
Keywords: Denoising, Noise Modeling, Poisson-Gaussian, Image Dataset, RAW, Smartphone, Calibration
@@ 1 Introduction
All digital images contain noise as a by-product of the capture process. Noise arises from the quantum effects of photon arrival, and from various electronic components in the sensor and camera circuitry, including sensor readout, amplifiers, and analog-to-digital conversion. Images captured in low-light environments, and especially with the smaller sensors found on smartphones, typically include considerable amounts of noise. This negatively affects the appearance and aesthetics of the images. Thus, noise reduction, or denoising, is an important and difficult task that has been the focus of active research for years. Denoising is an ill-posed inverse problem with no one correct solution, but rather an infinite set of possible solutions. Denoising is best performed on RAW images (rather than camera-generated JPEGs) because RAW images are scene-referred, undergo less processing by the image signal processor (ISP), contain linear data, and do not include gamma correction.
Currently, most research in denoising focuses on machine learning models. Thus, researchers developing new denoising algorithms rely on large image datasets to train and test their models. There are three main approaches to acquire an adequate dataset.
@@
  • Second, one can use an existing real noisy dataset that consists of low- and high-ISO images that have been extremely carefully aligned. Making a real noise dataset is exceptionally difficult, so there are only a few available. Most were created in 2017 or 2018, and thus use older cameras such as the iPhone 7, released more than nine years ago. Real noise datasets are also often limited to specially lit indoor scenes.
  • Third, one can add noise to one’s own images using a heteroscedastic Poisson-Gaussian noise model, which is superior to a homoscedastic one like AWGN. A heteroscedastic model can be easily retrieved from DNG images, though we found those models are often inaccurate and incomplete. We have chosen a variant of the third approach: create a custom physically-based heteroscedastic noise model by calibrating cameras and then refining the noise model with special tuning images. We use those models to introduce noise into clean reference RAW images. This approach requires more effort than simply using DNG noise profiles, but our models provide perfectly aligned, high-quality noisy RAW images. Further, this approach offers full control over the choice of cameras and scenes used.
The choice and quality of the training dataset directly affects the performance of a denoising model on real images. For physically modeled datasets, the key is minimizing the discrepancy between the synthetic noise and actual noise (Zhang et al., 2021). In this paper, we focus on physics-based noise modeling, using careful calibration to determine photon shot noise, combined with complete capture of signal-independent noise sources via dark frames. We confirm the inaccuracy of manufacturer DNG noise profiles. We demonstrate how to create and tune a heteroscedastic, physically-based noise model to produce accurate synthesized noise, and use our models to create a large synthesized noisy image dataset, with both RAW and TIFF images. Our noisy images produce excellent results both in LPIPS measurements and in denoising tests.
@@
  • Heteroscedastic noise models that achieve a 54% to 64% reduction in PSNR gap vs. standard DNG noise models.
@@ 2 Related Work
This section discusses the principal ways in which noise models and datasets are created to train machine learning denoising models. While learning-based noise generation methods exist (Abdelrahman et al., 2019), this paper focuses on physics-based statistical approaches for dataset creation, and we do not discuss denoising models in depth, of which there are many.
@@ 2.1 Additive White Gaussian Noise
Additive White Gaussian Noise (AWGN) is a simple noise model. AWGN adds noise to all pixels, regardless of intensity, using a simple Gaussian distribution with zero mean and a chosen variance (Abdelrahman et al., 2019):
@@ 2.2 Poisson-Gaussian Noise (P-G)
Poisson-Gaussian (P-G) is a physics-based noise model which improves upon AWGN by including both a signal-dependent component (shot noise), and a signal-independent component (read noise). Shot noise occurs because photon arrival is a quantum random process. For a pixel with mean signal $\mu$, the actual observed photon count includes random photon shot noise $N_{p}$. The total photon count follows a Poisson distribution (Zhang et al., 2021):
@@ 2.2.1 DNG Noise Profiles
Adobe, Inc. is the creator of the Digital Negative Specification (DNG) which is an open format for encoding digital images (Adobe Systems Incorporated, 2023). DNG files can store either RAW or non-RAW images. The DNG NoiseProfile tag (51041) uses the Poisson-Gaussian formulation. There are some weaknesses in the NoiseProfile specification. First, the specification simply contains one or more pairs of numbers (the scaling and bias parameters from Poisson-Gaussian). It does not include the ISO for which the noise applies, and there is no way to provide the parameters for multiple ISO values. Second, while some DNG NoiseProfiles include the parameters for each of the Bayer pattern channels (R, G1, G2, and B), it is very common to find NoiseProfiles that contain only three pairs of numbers (R, G, B), or simply a single pair of numbers which are then applied to all four channels equally (iPhone DNGs contain only this single pair). Since the ISO is not specified in the NoiseProfile tag, one must examine DNG files at different ISOs to determine the actual DNG model.
Finally, beyond the lack of per-channel information, DNG noise profiles provided by camera manufacturers tend to be rather inaccurate, compared to calibrated profiles (Zhang et al., 2021). We have found that profiles generated by Adobe for proprietary RAW files (such as the Sony ARW format) can be more accurate, but still retain the other issues described above.
@@ 2.2.2 Calibration and Dark Current
Modeling noise accurately using a Poisson-Gaussian formulation requires careful calibration (described in Methods). Healey and Kondepudy (1994) discuss calibrating a CCD using a gray card, as well as extracting Fixed Pattern Noise by shifting both the light source and gray card. In our calibrations, we confirmed the finding by Zhang et al. (2021) that fixed pattern noise is not a relevant component in modern CMOS sensors. Thus, our method for calibration deviates somewhat from Healey and Kondepudy (1994) and includes a tuning portion using real noisy images at a range of ISO levels.
The signal-independent part of noise appears in all images from a given sensor, even completely dark ones (e.g., with a lens cap on). Any electronics component can produce noise (Zhang et al., 2021); as a result, it is extremely difficult, if not impossible, to fully characterize and model this noise physically. Zhang et al. (2021) provide the complete formulation for a digital image D:
@@
where $K_{a}$ and $K_{d}$ represent the analog and digital gains, $N_{1}$ includes noise affected by analog gain, $N_{2}$ includes thermal and column fixed pattern noise that takes place after the analog gain step, and $N_{q}$ is quantization noise when converting analog values to digital ones.
There are a few techniques for capturing signal-independent noise for the purpose of noise synthesis. The simplest (and most common) is to compute it as part of the linear regression required for P-G noise modeling. The second is to capture dark frames (frames with no light hitting the sensor, such as with a lens cap, and/or in a completely dark room). This second approach, described by Zhang et al. (2021) as dark current, is used in lieu of the Poisson-Gaussian bias term. It more accurately captures sensor-independent noise, as dark images include all noise sources by definition. However, as different sensors (and camera electronics) produce different amounts of noise (and at different stages), we have found that the value of capturing dark current varies substantially from sensor to sensor. Specifically, professional-quality cameras can be very accurately modeled using a standard P-G linear regression. Further, some image signal processors (ISPs) in cameras, such as the iPhone 11 Pro, will process RAW images to suppress or alter noise based on the contents of the scene.
@@ 2.3 Denoising Datasets
There are a number of published denoising datasets, falling into two categories: ones that use AWGN to synthesize noise, and ones that use real noisy images. Due to the quality issues with AWGN, and the lack of published P-G datasets, we only list real noise datasets. Table 1 lists some popular real noisy image datasets (Chen et al., 2018; Abdelhamed et al., 2018; Anaya and Ortiz, 2018; Plötz and Roth, 2017; Wei et al., 2020; Flepp et al., 2024; Xu et al., 2018).
@@
Table 1: Real Noise Datasets. Note: Most RAW datasets listed are from 2017-2018, highlighting a need for updated datasets.
Real noise datasets group low-ISO clean images with noisy images captured at various ISO values (Wei et al., 2020). Perfect alignment is challenging; even sub-millimeter camera movement can cause misalignment. Further, strict lighting requirements limit most datasets to indoor scenes. For example, SIDD contains 30,000 images but only 10 distinct scenes (Wei et al., 2020), while DND provides just 50 image pairs (Plötz and Roth, 2017). Most real noise datasets were created in 2017-2018, predating recent sensor advances. Perfectly aligned real noise is a high bar, but when achieved, it provides an unimpeachable ground truth for training. Some datasets, including MIDD (Flepp et al., 2024) and PolyU (Xu et al., 2018), only include JPEG or PNG images, so they are not usable for RAW-based denoising.
Aside from AWGN, the alternative to real noise is physically modeled, Poisson-Gaussian synthesis. DNG RAW images include P-G noise profiles provided by the manufacturer or by Adobe as part of its camera calibration process. Because the quality of P-G models varies greatly, we created our own high-quality noise models. While somewhat laborious to create, our models have advantages over real noise. Our noisy images are perfectly aligned by definition, both indoor and outdoor scenes can be captured with a variety of lighting, and there is no limit to the number of images that can be produced. Further, by using dark current frames when appropriate, complex noise sources can be modeled.
@@ 3.1 RAW Images and Noise
Most cameras (including the ones we used) use a CMOS chip with a color filter array placed over the sensor sites. The color filter array usually follows the Bayer pattern (also called a mosaic), which arranges red, green, and blue filters in a prescribed order. There are twice as many green pixels as either red or blue pixels, due to the human eye’s increased sensitivity to green. There are two principal ways to process images for noise modeling and synthesis. One is to demosaic the Bayer data into RGB data and measure that data (or simply use pre-rendered TIFF, PNG, or JPEG data). The second is to operate on the Bayer data directly. Our approach works with the Bayer data directly because it provides higher quality. Bayer data is linear, rather than gamma-corrected, and we can ignore white balance and the specifics of the demosaic algorithm or the tone curve applied (of which there are many). Further, the RAW-based noise model can produce different values for each of the two green channels which is not possible once they have been processed into RGB.
@@ 3.2 DNG Noise Profile Analysis
The Poisson-Gaussian formulation enables straightforward synthesis of noisy images. However, the quality of this synthesis is entirely dependent on the accuracy and completeness of the formulation. One can either build a noise model manually, or rely on published noise profiles. Because building profiles manually can be labor-intensive and error-prone, most researchers rely on profiles stored as DNG NoiseProfile tags. For proprietary RAW files, the simplest approach is to convert RAW files into DNGs and then read the NoiseProfile tag written by Adobe’s DNG Converter software. This noise profile is computed by Adobe as part of its process for supporting proprietary RAW files. As mentioned previously, the DNG NoiseProfile specification does not include ISO data, so to build an accurate P-G model, one must provide the DNG converter a series of RAW files, shot with the desired range of ISO values. We found that for two Sony cameras (RX100 IV and A7R III), the Adobe profiles are very similar to the ones we built through our calibration and include noise parameters for red, green, and blue channels (though not for both green channels).
For cameras (including smartphone cameras) that produce DNG files natively, the noise profiles are provided by the manufacturer, not Adobe. We found that the DNG noise profile for the iPhone 11 Pro is both incomplete and inaccurate, especially at higher ISO values, which confirms the findings of Zhang et al. (2021). First, the DNG NoiseProfile for the iPhone 11 Pro does not include scaling and offset values for each channel, providing instead only a single pair of values. Second, as the ISO rises, the scaling factor should rise, since noise is related to analog gain (ISO). However, for the iPhone 11 Pro, the values level off. This is discussed further in the next section.
@@
Figure 3: DNG NoiseProfile scaling term (a) for iPhone 11 Pro and iPhone 15 Pro across ISO values. Both models show unexpected non-monotonic behavior at high ISOs.
@@ 3.3 Non-linearity in iPhone Noise Models
We observed a distinct difference between the noise models for the Sony cameras and the iPhone main and telephoto cameras. While the Sonys had a fairly consistent linear relationship (in log-log space) with ISO, the models for the iPhone sensors sharply level off, which is physically impossible. The Sony noise models were created by Adobe, while the iPhone captures in DNG format already, so the models are written into the original files by Apple. This leveling-off would point to one of two explanations: either the calibration was performed incorrectly, or the images used to compute the noise parameters do not show the expected noise. Because we replicated this same leveling-off in our own calibration process for the iPhone cameras, it pointed to the latter explanation. Visually inspecting RAW DNG images confirmed our suspicions: at the same ISO and exposure time, despite being Bayer pattern RAW images, our flat-field iPhone images do not have the same noise characteristics as real noisy RAW images containing texture and complexity.
@@
Tuning Image
Figure 4: Crops of calibration and tuning iPhone 11 Pro images shot at the same ISO and exposure times.
Thus, to produce iPhone images with accurately synthesized noise, we found our calibrated model must be further adjusted using carefully captured ”tuning images” at a range of ISO values.
We tuned and evaluated our models using 6 scenes captured across the four cameras (3 scenes each captured with the iPhone and Sony RX100 IV, and 3 additional scenes captured with the Sony A7R III). Each scene was captured at a low ISO (clean reference) and at 6-7 noisy ISO levels in powers of two, with two noisy captures per ISO level for baseline noise analysis. This yielded 156 full-resolution RAW tuning images spanning most or all of the true ISO range of each camera. All tuning images were captured using the same tripod and tethering setup as calibration to minimize motion (described in Calibration Procedure). Unlike the flat-field calibration images, the tuning images contained diverse textures, colors, and content, with a mix of indoor and outdoor scenes.